• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 19:57:07 on January 10, 2009

    Entry number 255360

    keyword=BB Cerenkov noise tests

    Run 4848:
      - HVs for 1--10 on, pulser on
      - LED #5, #10 plugged in
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      = Noise present on ADC inputs 1, 2 (2nd and 3rd)
    
    Run 4849:
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      - swapped 60Hz filter at ADC:  WAS #6, now a new filter "#12"
      = No effect on noise at input 1, 2
    
    Run 4850:
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      - swapped signal inputs 0, 1 at the patch panel just upstream of
        the ADC
      = Noise followed the cable
    
    Run 4851:
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      - swapped signals back to original inputs
      - HV off on 2nd PMT (ADC input 1)
      = Noise is still there (unaffected by HV on/off)
    
    Run 4852:
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      - HV still off on 2nd PMT (ADC input 1)
      - swapped signal cables right at the PMTs for PMT 1, 2
      = Noise unchanged (i.e. NOT coming from the PMT/base)
    
    Run 4853:
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      - HV still off on 2nd PMT (ADC input 1)
      - signal cables swapped back at the PMTs (now normal)
      - swapped signal cables right at the front-end electronics
        amplifier inputs for PMTs 1, 2
      = Noise stayed on same ADC channels, but the right-edge amplitude
      went from channel 400 to 600...  The little peak around channel
      110 also seems suppressed now.
      = ADC channel 0 has _no_ noise in it, not even the low-level
      'normal' background.  It is now viewing PMT 2 that has the HV
      turned off.
    
    Run 4854:
      - same setup as 4853, just jiggled the cables a bit before
        rerunning to make sure they're seated.
      = same results as 4853
    
    Run 4855:
      - same setup as 4853, but HV OFF on PMT 1 now (as well as off on
        PMT 2)
      = No effect
    
    Run 4856:
      - ALL HVs turned OFF, signals 1,2 still swapped at FE amplifier.
      = No effect on noisy ADC inputs 1, 2, but no background noise in
      ADC 3 channel any more (it goes away when the associated PMT HV is
      Off).
      = It looks like the problem is downstream of the FE amplifier, and
      upstream of the ADC.
    
    [ Fed pulser into PMT 2 signal cable at PMT to confirm where the
      signal came out at the DAQ weldment, Pulser stayed unplugged after
      this. ]
    
    Run 4857:
      - ALL HVs turned OFF, 
      - signals 1,2 swapped back at FE amplifier (normal now)
      - swapped the PMTs signals at the first patch panel at the DAQ
        weldment (ie. cables E-01R and E-02R).
      = noise was on ADC 0, not on ADC 1
    
    Run 4858:
      - ALL HVs turned OFF, 
      - swapped back the PMTs signals at the first patch panel at the
        DAQ weldment (all back to normal now)
      = ALL noise went away now... everything looks good... weird.
    
    Run 4859:
      - ALL HVs turned OFF, 
      - swapped signal cables 1,2 at the the first patch panel at the
        DAQ weldment again (same as run 4857)
      = Everything is still clean.   Problem solved itself...  Bad
      cable, flaky connection, who knows...
    
    [ Plugged pulser back into PMT 20's LED input.  This is the
      configuration from run 4857. ]
    
    Run 4860:
      - ALL HVs turned OFF, 
      - pulser plugged back in to LED input on PMT 20 (now the same as
        for run 4857)
      - signal cables 1,2 still swapped at the the first patch panel at
        the DAQ weldment (same as run 4857)
      = noise is back on ADC inputs 0 and 2
    
    Run 4861:
      - ALL HVs turned OFF, 
      - signal cables 1,2 swapped back at the the first patch panel at
        the DAQ weldment
      - all signal cables back where they should be
      = all noise is gone again...
    
    Run 4862:
      - repeat of 4860 configuration
        - cables 1,2 swapped at DAQ weldment PP
      = noise is back on ADC inputs 0 and 2
    
    Run 4863:
      - repeat of run 4859
      - pulser unplugged from PMT 20 LED input again
      = noise remains (contrary to run 4859)
    
    Run 4864:
      - cables 1,2 swapped back at DAQ weldment
        - all signal cables back to normal again.
      - this is the same as 4861, but with pulser unplugged from PMT 20
        LED input
      = all noise is gone again...
    
    Run 4865:
      - second run with 4864 config (no changes at all)
      = still clean (no noise)
    
    Run 4866:
      - turned on HV for PMT 2 only
      = noise back on ADC channels 1 and 2
    

    Data are inconsistent... Not sure where the problem lies. Will try again tomorrow.